OpenStudy (ariane12):

I'll medal and fan.. please help.. find the equation of the bisector of the interior angles of the triangle bounded by the lines x=3, y=4 and x+y=2.

OpenStudy (boldjon):

i got u

OpenStudy (boldjon):

the slopes of the lines in question are 0, undefined and 1 this is a 45-45-90 right triangle. the vertexes are at (3,4), (3,-1), (-2,4) L1 through (3,4) and point of intersection. this one we can jump to the chase... we know that it bisects the right angle, it is 45 degrees... the slope is 1. (y-4) = (x-3) y = x + 1 L2 through (-2,4) the slope... is a 22.5 degree angle and it is downward sloping... - 22.5 degrees tan -22.5 =- sqrt ((1-cos 45)/(1+cos45)) = 1-sqrt 2 (y-4) = (1-sqrt2) (x+2) L3 through (3,-1) is downward sloping and is a 67.5 degree angle. =- sqrt ((1+cos 45)/(1-cos45)) = -1-sqrt 2 (y+1) = (-1-sqrt2) (x-3)